OpenStudy (anonymous):

At a stoplight, a truck traveling at 15 m/s passes a car as it starts from rest.?

OpenStudy (anonymous):

i dont know how to combine the equations

OpenStudy (anonymous):

what u wanna find ???

OpenStudy (anonymous):

At a stoplight, a truck traveling at 15 m/s passes a car as it starts from rest. The truck travels at constant velocity and the car accelerates at 3 m/s^2. How much time does the car take to catch up to the truck?

OpenStudy (anonymous):

I want to know how to solve it step by step :\

OpenStudy (anonymous):

i know that some how we get to this but i dont understand how, 15t=1/23t^2

OpenStudy (anonymous):

15t=(1/2)3t^2

OpenStudy (anonymous):

so find t

OpenStudy (anonymous):

ohh srry

OpenStudy (anonymous):

i dont know how to get that though? i copied it from the board the teacher didnt explain it

OpenStudy (anonymous):

u see car moves with constant velocity . u let t be d time elapsed in which d whole thing takes place n as we can see that the displacement of car {(1/2)3t^2} n truck(15t) will be equal

OpenStudy (anonymous):

therefore (1/2)3t^2= 15t

OpenStudy (anonymous):

is there a formula that you used to substitute the initial velocity and acceleration

OpenStudy (anonymous):

see when there is constant acceleration then we have 3 eqns v = u +at s=ut+1/2at^2 2as=v^2-u^2 where u is initial velocity , v is final velocity s is displacement n a is cnstant acceleration

OpenStudy (anonymous):

i used d 2nd one 4 finding displacement of car in time t

OpenStudy (anonymous):

hello is it clear ?????

OpenStudy (anonymous):

hey sorry

OpenStudy (anonymous):

hey dude r u satisfied with d ans n sorry stands 4 what??????

OpenStudy (anonymous):

oh i was away yeah ill figure it out, my teacher uses different symbols fir final velocity and initial and displacement

OpenStudy (anonymous):

kk fine u just put d values n symbols do not matter anyways ,,,,,,,, bbyeeeee

OpenStudy (anonymous):

Well explain

OpenStudy (anonymous):

The car will overtake the truck when it catches up. The technical way of saying that is that they will both have gone the same distance from the traffic lights. That's it. that's what overtaking means. The distance one goes under constant acceleration is described by the equation given above by arzoo: s (distance) = u(initial velocity)*t(time) + 1.2 * a(acceleration) * t(time)^2 The truck doesn't accelerate but just goes at a constant speed. So its a is zero and its distance s is s(truck) = u*t = 15t The car starts at u=zero but does accelerate. So its distance is s(car)= 1/2 * a * t^2 = 1/2 * 3 * t^2 And these are equal when the car overtakes - because it has caught up and the distances are the same. So... s(truck)=s(car) 15t = 1/2 * 3 * t^2 Rearrange and cancel out... Just as arzoo says.